Circular No. 5945

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

NOVA SAGITTARII 1994 M. Della Valle, Padua; P. Molaro, Osservatorio Astronomico di Trieste;

J. Casares, Astrophysics Department, Oxford University; and E. Martin, Instituto de Astrofisica de Canarias, Tenerife, report: “Spectrograms (range 380–860 nm, resolution ∼ 3 nm) obtained on Mar. 7.2 UT with the New Technology Telescope (+ EMMI) at La Silla suggest this object to be a nova caught during its early decline. The spectrum shows Hα and Na I in emission, flanked by blueshifted P-Cyg profiles, superimposed onto a strong continuum. The mean expansion velocity derived from the minimum of the P-Cyg absorptions is ∼ 2200 km/s.”

A0535+26 R. B. Wilson, B. A. Harmon, and G. J. Fishman, Marshall Space Flight

Center, NASA; and M. H. Finger, Universities Space Research Association, report for the Compton Gamma Ray Observatory BATSE team: “The strong outburst from A0535+26 (IAUC 5931, 5933, 5934) peaked on Feb. 18, reaching a total intensity of ∼ 8 Crab in the band 20–40 keV. The phase-averaged pulsed flux then was 1.8 Crab (total). The intensity has declined approximately linearly since then, reaching ∼ 4 Crab (total) and 1.1 Crab (pulsed) on Feb. 26. During the peak of its outburst the pulse frequency was increasing at a rate of ∼ 1.2 × 10^{−11} s^{−2} or P/Ṗ ∼ 25 yr.”

NOVA CASSIOPEIAE 1993 Photoelectric photometry by M. Winiarski and M. Kurpinska-Winiar-

ska, Cracow Astronomical Observatory (W; comparison star SAO 35665: V = 7.24, B − V = +0.17) and H. Mikuž, Ljubljana, Slovenia (M; IAUC 5929): Feb. 13.784 UT, V = 9.16 ± 0.02, B − V = +0.24 ± 0.02 (W); 14.766, 9.31 ± 0.01, +0.25 ± 0.01 (W); 15.739, 9.82 ± 0.01, +0.28 ± 0.01 (W); 16.242, 10.07 ± 0.02, +0.26 ± 0.02 (W); 17.279, 10.41 ± 0.05, +0.34 ± 0.06 (W); 25.83, 14.27 ± 0.04, – (M); Mar. 4.80, 15.03 ± 0.05, – (M).

PERIODIC COMET SCHWASSMANN-WACHMANN 1 Total visual magnitude estimates by L. Szentaskó at Veresegyház, Hun-

gary (0.33-m reflector): Feb. 16.79 UT, 13.0; Mar. 3.81, 12.3; 5.79, 13.0.

1994 March 7 Daniel W. E. Green